OpenStudy (graciebugg):

HELP Will Fan and Medal... Thanks in Advance <8th Grade Math> Which shows the expression below in simplified form? (6.4 × 10^-4) - (3.8 × 10^-5) A. 2.6 × 10^-4 B. 2.6 × 10^-5 C. 6.02 × 10^-4 D. 6.02 × 10^-5

OpenStudy (graciebugg):

@hysusonic

OpenStudy (graciebugg):

@barakahcharm

OpenStudy (graciebugg):

I figured it was either A or B. If you subtract 6.4 and 3.8 you get 2.6. The only thing that gets me is what exponent do I put?

OpenStudy (graciebugg):

@Willie579

OpenStudy (willie579):

(6.4 × 10-4) - (3.8 × 10-5) = 27.

OpenStudy (lochana):

\[6.4 × 10^{-4} - 3.8 × 10^{-5} = 6.4 × 10^{-4} - 0.38 × 10^{-4}\]

OpenStudy (willie579):

Well, I was trying to figure it out. :P

OpenStudy (lochana):

\[6.4 × 10^{-4} - 0.38 × 10^{-4} = 6.02 × 10^{-4}\]

OpenStudy (lochana):

try another one:)

OpenStudy (graciebugg):

@lochana can you explain that? I have another one just like it and I have no clue how to do it.

OpenStudy (graciebugg):

Also I had to make an edit to the question. ^ means the following number is an Exponent.

OpenStudy (lochana):

yes. if we have things like \[a.10^b \ , b.10^b\] you can sum them up such that \[(a+b).10^b\]

OpenStudy (lochana):

notice exponent of 10 should be same to do that.

OpenStudy (graciebugg):

Ok. Can you help me with the second one? I'll make another thread if so.

OpenStudy (lochana):

sure

OpenStudy (anonymous):

wait to the question is answered?